# For Fine-Tuned Models

### v1

@@ 1.1 Contributions
(a) Last-layer output changes on one batch of train data.
@@ 2 Related work
Delta-parameter pruning. DPs represent the parameter changes induced by fine-tuning and are critical to various operations. In Federated and Distributed Learning, DPs are continuously exchanged between a central server and distributed clients during training (Li et al., 2021, 2020; Khirirat et al., 2018; Chen et al., 2021). In model merging, the goal is to merge multiple task-specific models, fine-tuned from the same pretrained backbone, by combining their DPs into a single model with diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022). Similarly, in multi-task serving, DPs are processed separately while the base model weights are computed once during the forward pass (Yao & Klimovic, 2023; Liu et al., 2024). (Yu et al., 2023a; Liu et al., 2024; Yao & Klimovic, 2023) focus on reducing the size of DPs, and recently(Yu et al., 2023a) has popularized DARE as a poerful random-based DPP method. However, DARE fails with large DPs or high pruning rates. We propose techniques that, applied during both fine-tuning and post-hoc stages, restore performance and significantly improve upon vanilla DARE. Our results demonstrate, for the first time, that DPP remains effective even at extreme pruning rates, reducing up to 99% of the model parameters. Thanks to our modification, we also show that random-based DPP can be combined with LoRA and can be used for structured pruning.
@@ 3.2 Randomized DPP: Random Drop and Rescale (DARE)
Thus, the key factors influencing the magnitude of $|h_{i}^{\rm{diff}}|$ are: (i) the rescaling factor $1/(1-p)$ and (ii) the mean and variance (averages are over input dimension $j\in[n]$) of the influence parameters $\{c_{ij}\}$. Specifically, increasing the rescaling factor (eqv. increase drop-rate $p$) increases $|h_{i}^{\rm{diff}}|$ at a rate of $\mathcal{O}((1-p)^{-\frac{1}{2}})$, which can be large in demanding pruning scenarios when $p$ is large. Also DARE yields smaller $|h_{i}^{\rm{diff}}|$ values when $c_{ij}=\Delta W_{ij}x_{j}$ exhibit low first (mean) and second order (variance) averages. We validate these conclusions through elaborate experiments on a controlled setup in Appendix  B. More importantly, our algorithms in Sec. 5, further tested on LLMs, are also built upon the insights from this theorem and its observations.
@@ 4.1 Adjusting the rescaling factor
Motivation. Recall from Thm. 3.1 as $p$ increases, the absolute difference in outputs $|h_{i}^{\rm{diff}}|$ grows at a rate of $\mathcal{O}((1-p)^{-\frac{1}{2}})$. Setting the rescaling factor to $q=1-p$ eliminates the mean of $h_{i}^{\rm{diff}}$, but does not minimize its magnitude. Through empirical analysis, we demonstrate in Fig. 1 how various values of $q$ influence model outputs in terms of both mean output change (i.e., the average of $|h_{i}^{\rm{diff}}|$ across all last-layer neurons) in Fig. 1(a), and, test performance in Fig. 1(b) across different datasets. For 99% pruning rate ($0.011-p=$ 0.01) we observe these quantities over a range of $q$ values from $0.005$ to $0.04$. The results depicted in Fig. 1(a) for mean output change, indicate a convex-like trend, with minimum points (marked with asterisks) achieved at $q$ values higher than the vanilla $1-p$. E.g., the optimal $q$ for COLA and SST2 is $\approx 0.018$ and $0.033$, respectively. Recall now that in our analysis, the amount of change in output activations is used as a proxy for test performance: smaller mean output change corresponds to better performance. Fig. 1(b) validates this: the test-performance curves show a concave-like shape, with the peak performance also occurring at $q>1-p$ and numerically consistent with the values of that minimize the mean absolute output change.
@@
DAREx-q with per-layer tuning. The implementations of DAREx-q described above select a single, global rescaling factor ($1/q_{v}$ or $1/q_{e}$) that is then applied to all surviving DPs across all layers of the network. However, since the DPs in intermediate layers each have their own mean output change over their respective output layers, a natural question arises: Can we improve performance by selecting a different rescaling factor for each layer $\ell\in[L]$? Moreover, can this be done efficiently without requiring $L$ hyperparameter searches (one for each layer), which would be computationally prohibitive? In Appendix C.8, we demonstrate that the answer to both of these questions is affirmative. To circumvent the computational challenge of $L$ searches, we leverage theoretical insights by extending Thm. 3.1 to provide a high-probability bound for arbitrary $q$ and then optimizing over it. Like Thm. 3.1, the new bound still depends on the first- and second-order statistics of the coefficients $c^{\ell}_{ij}$, which means the optimal $q^{\ell}$ can vary per layer $\ell$ based on these statistics. We denote this vector of per-layer rescaling factors as $1/\mathbf{q}_{v}=[1/q_{v}^{1},\ldots,1/q_{v}^{L}]$ and $1/\mathbf{q}_{e}$, respectively, for two implementations depending on whether the performance on validation set or mean output change is optimized. Please refer to App. C.8 for details. In Sec. 5.1, we show that this approach performs very well on encoder models and is comparable to the global $1/q$ approach for decoder models.
@@ 5 Experiments
Datasets and models for Encoder-based and Decoder-based LMs. For Encoder-based LMs, we utilize four datasets—sentence acceptability dataset CoLA (Warstadt et al., 2019), sentiment detection dataset SST-2 (Socher et al., 2013), paraphrase dataset MRPC (Dolan & Brockett, 2005), and sentence similarity dataset STS-B (Cer et al., 2017). For the selection of pretrained backbones, we choose BERT-base-uncased (Devlin et al., 2018) and RoBERTa-base (Liu et al., 2019), fine-tuning them on these task-specific datasets to obtain supervised finetuned models. For Decoder-based LMs, we focus on mathematical reasoning tasks. Due to constraints on resources for fine-tuning, we opted to fine-tune a smaller Qwen2-0.5B (Yang et al., 2024) model on the MetaMath dataset (Yu et al., 2023b). Additionally, we utilized publicly available mathematical reasoning models, including the MetaMath-llema-7B (Yu et al., 2023b), MetaMath-7B (Yu et al., 2023b), WizardMath-7B (Luo et al., 2023) and Abel-7B (Chern et al., 2023), all based on the Llama2-7B architecture (Touvron et al., 2023). We then use the GSM8K (Cobbe et al., 2021) to test these models.
Evaluation metrics. Following (Yu et al., 2023a), we evaluate performance using the Matthews correlation coefficient for CoLA, accuracy for SST-2, the average of accuracy and F1 score for MRPC, the average of Pearson and Spearman correlation for STS-B, and zero-shot accuracy for GSM8K.
@@ 5.1 Rescaling-parameter modification
We evaluate our DAREx-q algorithm. From Sec. 4.1, recall the four proposed variations (i.e., $1/q_{v},1/q_{e},1/\mathbf{q}_{v}$ and $1/\mathbf{q}_{e}$) for choosing the rescaling factor along two orthogonal axes: (a) whether to use labeled validation data or not ($1/q_{v}$ or $1/q_{e}$, respectively); and (b) global tuning versus per-layer tuning ($1/q$ vs $1/\mathbf{q}$).
@@
Encoder models. Table 2 compares DAREx-q (all four variants) with vanilla DARE (with rescaling factor $1/(1-p))$. We report average performance and standard deviation on the test set over four independent runs. For reference, we also report performance of the finetuned model before pruning. We find that: (1) All DAREx-q variants outperform vanilla DARE in all cases and by a large margin (e.g., $>40\%$ on COLA). (2) $1/q_{e}$ and $1/\mathbf{q}_{e}$ deliver competitive performance compared to $1/q_{v}$ and $1/\mathbf{q}_{v}$, making it an effective alternative when a labeled validation set is unavailable. (3) Per-layer rescaling ($1/\mathbf{q}_{v}$ and $1/\mathbf{q}_{e}$) outperforms global rescaling.
@@ 5.2 AdamR-$L_{2}$ finetuning
Encoder models. Table 5 compares the performance of DARE applied to the original finetuned model against its performance when applied to models finetuned with AdamR-$L_{2}$ across two encoder models and four datasets. We set the pruning rate at $p=0.99$ and, for reference, also report the performance of the finetuned models without pruning, both with and without AdamR-$L_{2}$. Consistent with the insights of Sec. 4.2, applying DARE after AdamR-$L_{2}$ (DAREx-$L_{2}$) outperforms vanilla DARE without regularization. Additionally, the performance is only marginally lower than that of the original finetuned model. Figure 7 in the Appendix further shows that increasing the regularization strength allows pruning up to 99.9% of parameters with only a minor performance drop. See Appendix C.3 for further ablations on regularization weight. Finally, Table 9 in the Appendix demonstrates that DAREx-$L_{2}$ unlocks the potential of random-based DPP for structural pruning, improving DARE without $L_{2}$ finetuning by over 40%.
@@ 5.3 How to apply importance-based DPP and when is it competitive?
Improvements with AdamR-$L_{1}$. For completeness of our study, we also demonstrate that finetuning with AdamR-$L_{1}$ can boost performance of importance-based DPP as suggested in Sec. 4.2. Specifically, in Table 7 in App. C.4, we show elaborate experiments comparing the performance of MP and WANDA (applied to DPs) when used on a model finetuned with/without AdamR-$L_{1}$ for varying values of pruning rates and two encoder models across four datasets. In summary, we find: (1) AdamR-$L_{1}$ consistently increases performance of MP. (2) While AdamR-$L_{1}$ improves WANDA’s performance for most datasets, this is not always the case. We attribute this to WANDA’s importance score being impacted not only by DPs $\Delta W_{ij}$ but also by input magnitude $|x_{j}|$ and important DPs not always being aligned with outlier activations (see App. D).
@@ 6 Conclusions and limitations
Our systematic study of DPP methods, provides theoretical insights, extensive experiments in controlled environments, and demonstrates practical applications on large-scale models and datasets. Starting with an in-depth analysis of random-based DPP methods, we identify key factors that degrade their performance. Based on these insights, we propose: (i) DAREx-q, a post-training rescaling modification strategy that significantly outperforms the state-of-the-art at high pruning rates and performs at least as well across other rates; and (ii) AdamR finetuning to enhance existing models that struggle with DPP, ensuring the production of highly prunable DPs. Our results, including dramatic gains over baseline methods shown in Table 1, suggest that these strategies warrant further investigation across different tasks and modalities. Additionally, preliminary results suggest that DAREx-q can be effectively combined with parameter-efficient finetuning methods like LoRA and utilized for structural DPP, both of which highlight the potential for further research. Finally, while finetuning LLMs typically results in small DPs favoring random-based DPP, we demonstrate that importance-based DPP methods can serve as robust alternatives for larger DPs. This highlights the need for further exploration of alternatives. In conclusion, our comprehensive study of DPP methods provides a framework for selecting appropriate DPP methods under different practical scenarios, which we summarize in Fig. 4.
Framework for DPP method selection: We outline the procedure for selecting appropriate DPP methods in practice. As illustrated in Fig 4, if fine-tuning is not permitted and the data points (DPs) have large statistics, we recommend using WANDA for DPP (see Sections 5.3). If the DPs are not large, DAREx-q with $1/q_{v}$ should be applied when a validation set is available, otherwise, DAREx-q with $1/q_{e}$ is recommended (see Sec. 4.1). If the existing DPs are insufficient and fine-tuning is allowed, we suggest using AdamR-$L_{2}$ or $L_{1}$ with appropriate regularization weights to generate highly prunable DPs for DAREx-q and MP, respectively (see Sec. C.6). Among the two, AdamR-$L_{2}$+DAREx-q (DAREx-$L_{2}$-q) should be prefered due to flexibility and better performance as shown in Table 9 and Table 5.
Broader impact: As discussed in Section 1, DPP offers potential advantages for online serving, gradient communication in Federated Learning, and storage saving. Our systematic study of DPP methods and the proposed strategies that improve performance of existing methods can positively impact the usage of LLMs. We also analyzed the impact of our approach on the efficiency of the machine learning system in App.  C.12.
Acknowledgments: This work was partially funded by an Alliance Mission Grant, the NSERC Discovery Grant No. 2021-03677, the Alliance Grant ALLRP 581098-22, NFRFE-2023-00936, Natural Sciences and Engineering Research Council of Canada (NSERC), Compute Canada, and Vector Institute.
@@ C.2 Implementation details
C.3 Controlling pruning with regularizaiton
@@ C.4 AdamR-$L_{1}$ finetuning
C.5 Scale of delta parameters
@@ C.9 Structural pruning
C.10 Comparison to sparse finetuning
@@
C.11 Random-based methods generally outperforms importance-based DPP
@@ Appendix E Proofs
For completeness, we visualize the three bounds that we discussed above in Figure 10. The lower the better, thus note how the Kearns-Saul bound improves over naive Hoeffding and the Berend-Kontorovich bound improves over both for $p>1/2$.

### v2

# Christos Thrampoulidis1

@@ 1.1 Contributions
(a) Last-layer output changes on a batch of training data.
@@ 2 Related work
Delta-parameter pruning. DPs represent the parameter changes induced by fine-tuning and are critical to various operations. In Federated and Distributed Learning, DPs are continuously exchanged between a central server and distributed clients during training (Li et al., 2021; 2020; Khirirat et al., 2018; Chen et al., 2021). In model merging, the goal is to merge multiple task-specific models, fine-tuned from the same pretrained backbone, by combining their DPs into a single model with diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022). Similarly, in multi-task serving, DPs are processed separately while the base model weights are computed once during the forward pass (Yao & Klimovic, 2023; Liu et al., 2024). (Yu et al., 2023a; Liu et al., 2024; Yao & Klimovic, 2023) focus on reducing the size of DPs, and recently(Yu et al., 2023a) has popularized DARE as a powerful random-based DPP method. However, DARE fails with large DPs or high pruning rates. We propose techniques that, applied during both fine-tuning and post-hoc stages, restore performance and significantly improve upon vanilla DARE. Our results demonstrate, for the first time, that DPP remains effective even at extreme pruning rates, reducing up to 99% of the model parameters. Thanks to our modification, we also show that random-based DPP can be combined with LoRA and can be used for structured pruning.
@@ 3.2 Randomized DPP: Random Drop and Rescale (DARE)
Thus, the key factors influencing the magnitude of $|h_{i}^{\rm{diff}}|$ are: (i) the rescaling factor $1/(1-p)$ and (ii) the mean and variance (averages are over input dimension $j\in[n]$) of the influence parameters $\{c_{ij}\}$. Specifically, increasing the rescaling factor (eqv. increase drop-rate $p$) increases $|h_{i}^{\rm{diff}}|$ at most a rate of $\mathcal{O}((1-p)^{-\frac{1}{2}})$, which can be large in demanding pruning scenarios when $p$ is large. Also DARE yields smaller $|h_{i}^{\rm{diff}}|$ values when $c_{ij}=\Delta W_{ij}x_{j}$ exhibit low first (mean) and second order (variance) averages. We validate these conclusions through elaborate experiments on a controlled setup in Appendix  B. More importantly, our algorithms in Sec. 5, further tested on LLMs, are also built upon the insights from this theorem and its observations.
@@ 4.1 Adjusting the rescaling factor
Motivation. Recall from Thm. 3.1 as $p$ increases, the absolute difference in outputs $|h_{i}^{\rm{diff}}|$ grows at at most [^3] The rate of growth with $p$ in the high-probability upper bound of Thm. 3.1 is tight and thus predictive of the output change taking larger values as $p$ increases; see App. E.1. a rate of $\mathcal{O}((1-p)^{-\frac{1}{2}})$. Setting the rescaling factor to $q=1-p$ eliminates the mean of $h_{i}^{\rm{diff}}$, but does not minimize its magnitude. Through empirical analysis, we demonstrate in Fig. 1 how various values of $q$ influence model outputs in terms of both mean output change (i.e., the average of $|h_{i}^{\rm{diff}}|$ across all last-layer neurons) in Fig. 1(a), and, test performance in Fig. 1(b) across different datasets. For 99% pruning rate ($0.011-p=$ 0.01) we observe these quantities over a range of $q$ values from $0.005$ to $0.04$. The results depicted in Fig. 1(a) for mean output change, indicate a convex-like trend, with minimum points (marked with asterisks) achieved at $q$ values higher than the vanilla $1-p$. E.g., the optimal $q$ for COLA and SST2 is $\approx 0.018$ and $0.033$, respectively. Recall now that in our analysis, the amount of change in output activations is used as a proxy for test performance: smaller mean output change corresponds to better performance. Fig. 1(b) validates this: the test-performance curves show a concave-like shape, with the peak performance also occurring at $q>1-p$ and numerically consistent with the values of that minimize the mean absolute output change.
@@
DAREx-q with per-layer tuning. The implementations of DAREx-q described above select a single, global rescaling factor ($1/q_{v}$ or $1/q_{e}$) that is then applied to all surviving DPs across all layers of the network. However, since the DPs in intermediate layers each have their own mean output change over their respective output layers, a natural question arises: Can we select a different rescaling factor for each layer $\ell\in[L]$? Moreover, can this be done efficiently without requiring $L$ hyperparameter searches (one for each layer), which would be computationally prohibitive? In Appendix C.9, we demonstrate that the answer to both of these questions is affirmative. To circumvent the computational challenge of $L$ searches, we leverage theoretical insights by extending Thm. 3.1 to provide a high-probability bound for arbitrary $q$ and then optimizing over it. Like Thm. 3.1, the new bound still depends on the first- and second-order statistics of the coefficients $c^{\ell}_{ij}$, which means the optimal $q^{\ell}$ can vary per layer $\ell$ based on these statistics. We denote this vector of per-layer rescaling factors as $1/\mathbf{q}_{v}=[1/q_{v}^{1},\ldots,1/q_{v}^{L}]$ and $1/\mathbf{q}_{e}$, respectively, for two implementations depending on whether the performance on validation set or mean output change is optimized. Please refer to App. C.9 for details. In Sec. 5.1, we show that this approach performs very well on encoder models and is comparable to the global $1/q$ approach for decoder models.
@@ 5 Experiments
Datasets and models for Encoder-based and Decoder-based LMs. For Encoder-based LMs, we utilize four datasets—sentence acceptability dataset CoLA (Warstadt et al., 2019), sentiment detection dataset SST-2 (Socher et al., 2013), paraphrase dataset MRPC (Dolan & Brockett, 2005), and sentence similarity dataset STS-B (Cer et al., 2017). For the selection of pretrained backbones, we choose BERT-base-uncased (Devlin et al., 2018) and RoBERTa-base (Liu et al., 2019), fine-tuning them on these task-specific datasets to obtain supervised finetuned models. For Decoder-based LMs, we focus on mathematical reasoning tasks. Due to constraints on resources for fine-tuning, we opted to fine-tune a smaller Qwen2-0.5B (Yang et al., 2024) model on the MetaMath dataset (Yu et al., 2023b). Additionally, we utilize publicly available mathematical reasoning models, including the MetaMath-llema-7B (Yu et al., 2023b), MetaMath-7B (Yu et al., 2023b), WizardMath-7B (Luo et al., 2023) and Abel-7B (Chern et al., 2023), all based on the Llama2-7B architecture (Touvron et al., 2023). We then use the GSM8K (Cobbe et al., 2021) to test these models.
Evaluation metrics. Following (Yu et al., 2023a), we evaluate performance using various metrics: the Matthews correlation coefficient for CoLA, accuracy for SST-2, a combined score of accuracy and F1 for MRPC, the mean of Pearson and Spearman correlations for STS-B, and zero-shot accuracy for GSM8K.
@@ 5.1 Rescaling-parameter modification
We evaluate our DARq algorithm, as discussed in Section 4.1. Recall the four proposed variations for choosing the rescaling factor, categorized by two distinct factors: (a) the use of labeled validation data or not ($1/q_{v}$ or $1/q_{e}$, respectively); and (b) global tuning versus per-layer tuning ($1/q$ versus $1/\mathbf{q}$).
@@
Encoder models. Table 2 compares DAREx-q (all four variants) with vanilla DARE (with rescaling factor $1/(1-p))$. We report average performance and standard deviation on the test set over four independent runs. For reference, we also report performance of the finetuned model before pruning. We find that: (1) All DAREx-q variants outperform vanilla DARE in all cases and by a large margin (e.g., $>40\%$ on COLA). (2) $1/q_{e}$ and $1/\mathbf{q}_{e}$ deliver competitive performance compared to $1/q_{v}$ and $1/\mathbf{q}_{v}$, making it an alternative when a labeled validation set is unavailable. (3) Per-layer rescaling ($1/\mathbf{q}_{v}$ and $1/\mathbf{q}_{e}$) outperforms global rescaling.
@@ 5.2 AdamR-$L_{2}$ finetuning
Encoder models. Table 5 compares the performance of DARE applied to the original finetuned model against its performance when applied to models finetuned with AdamR-$L_{2}$ across two encoder models and four datasets. We set the pruning rate at $p=0.99$ and, for reference, also report the performance of the finetuned models without pruning, both with and without AdamR-$L_{2}$. Consistent with the insights of Sec. 4.2, applying DARE after AdamR-$L_{2}$ (DAREx-$L_{2}$) outperforms vanilla DARE without regularization. Additionally, the performance is only marginally lower than that of the original finetuned model. Figure 7 in the Appendix further shows that increasing the regularization strength allows pruning up to 99.9% of parameters with only a minor performance drop. See Appendix C.4 for further ablations on regularization weight. Finally, Table 9 in the Appendix demonstrates that DAREx-$L_{2}$ unlocks the potential of random-based DPP for structural pruning, improving DARE without $L_{2}$ finetuning by over 40%. We also conduct training analysis of DAREx-$L_{2}$ in Appendix F.
@@ 5.3 How to apply importance-based DPP and when is it competitive?
Improvements with AdamR-$L_{1}$. For completeness of our study, we also demonstrate that finetuning with AdamR-$L_{1}$ can boost performance of importance-based DPP as suggested in Sec. 4.2. Specifically, in Table 7 in App. C.5, we show elaborate experiments comparing the performance of MP and WANDA (applied to DPs) when used on a model finetuned with/without AdamR-$L_{1}$ for varying values of pruning rates and two encoder models across four datasets. In summary, we find: (1) AdamR-$L_{1}$ consistently increases performance of MP. (2) While AdamR-$L_{1}$ improves WANDA’s performance for most datasets, this is not always the case. We attribute this to WANDA’s importance score being impacted not only by DPs $\Delta W_{ij}$ but also by input magnitude $|x_{j}|$ and important DPs not always being aligned with outlier activations (see App. D).
@@ 6 Conclusions and limitations
Our study of DPP methods, provides theoretical insights, extensive experiments in controlled environments, and demonstrates practical applications on large-scale models and datasets. Starting with an in-depth analysis of random-based DPP methods, we identify key factors that degrade their performance. Based on these insights, we propose: (i) DAREx-q, a post-training rescaling modification strategy that significantly outperforms the state-of-the-art at high pruning rates and performs at least as well across other rates; and (ii) AdamR finetuning to enhance existing models that struggle with DPP, ensuring the production of highly prunable DPs. Our results, including dramatic gains over baseline methods shown in Table 1, suggest that these strategies warrant further investigation across different tasks and modalities. Additionally, preliminary results suggest that DAREx-q can be effectively combined with parameter-efficient methods like LoRA and utilized for structural DPP, both of which highlight the potential for further study. Finally, while finetuning LLMs typically results in small DPs favoring random-based DPP, we demonstrate that importance-based DPP methods can serve as robust alternatives for larger DPs. This highlights the need for further exploration of alternatives. In conclusion, our comprehensive study of DPP methods provides a framework for selecting appropriate DPP methods under different scenarios, which we summarize in Fig. 4 (details in App. C.3). Broader impact see App.  C.13.
Acknowledgments: This work was partially funded by an Alliance Mission Grant, the NSERC Discovery Grant No. 2021-03677, the Alliance Grant ALLRP 581098-22, NFRFE-2023-00936, Natural Sciences and Engineering Research Council of Canada (NSERC), Canada CIFAR AI Chairs program, Canada Research Chair program, MITACS-CIFAR Catalyst Grant Program, the Digital Research Alliance of Canada, Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT).
@@ C.2 Implementation details
C.3 Proposed framework for DPP method selection.
We outline the procedure for selecting appropriate DPP methods in practice. As illustrated in Fig 4, if fine-tuning is not permitted and the data points (DPs) have large statistics, we recommend using WANDA for DPP (see Sections 5.3). If the DPs are not large, DAREx-q with $1/q_{v}$ should be applied when a validation set is available, otherwise, DAREx-q with $1/q_{e}$ is recommended (see Sec. 4.1). If the existing DPs are insufficient and fine-tuning is allowed, we suggest using AdamR-$L_{2}$ or $L_{1}$ with appropriate regularization weights to generate highly prunable DPs for DAREx-q and MP, respectively (see Sec. C.7). Among the two, AdamR-$L_{2}$+DAREx-q (DAREx-$L_{2}$-q) should be prefered due to flexibility and better performance as shown in Table 9 and Table 5.
C.4 Controlling pruning with regularizaiton
@@ C.5 AdamR-$L_{1}$ finetuning
C.6 Scale of delta parameters
@@ C.10 Structural pruning
C.11 Comparison to sparse finetuning
@@
C.12 Random-based methods generally outperforms importance-based DPP
@@ Appendix E Proofs
[…]
E.1.1 Lower bound